Two resistances 6 Ω and 3 Ω are connected in parallel. A current of 6 amperes is passing through the combination. Find:

a. The total resistance, and

b. The current through each resistor



1
Expert's answer
2021-05-31T02:08:06-0400

Total Resistance = 1/6 + 1/3

= 1+2/6 = 3/6

6/3 = 2

B) current through resistance

= 6/2 = 3 amps